# V2X-Radar: A Multi-modal Dataset with 4D Radar for Cooperative Perception

### v2

@@ 4.1 Single-agent 3D Object Detection
Evaluation Metrics. The evaluation area covers [-100, 100] in x direction and [0, 100] in y direction from the ego-vehicle or roadside unit. Following the metrics in Rope3D [46], KITTI [6] datasets, we utilize the Average Precision (AP) at IoU thresholds of 0.5 and 0.7 to evaluate the performance of object detection.
@@
Table 3: Roadside 3D object detection benchmarks on V2X-Radar-I under homogeneous split. The vehicle category includes car, bus and truck. ‘M‘ means modality, and L, C, and R denote LiDAR, Camera, and 4D Radar, respectively.
@@
Table 4: Single-vehicle 3D object detection benchmarks on V2X-Radar-V under homogeneous split. The vehicle category includes car, bus and truck. ‘M‘ means modality, and L, C, and R denote LiDAR, Camera, and 4D Radar, respectively.
@@
4.2 Cooperative 3D Object Detection
Evaluation Metrics. The evaluation area spans [-100, 100] metres in both the x and y directions relative to the ego vehicle. Similar to DAIR-V2X [48] and V2V4Real [39], we group various vehicle types into the same class and concentrate solely on vehicle detection. The performance of object detection is evaluated using Average Precision (AP) at IoU thresholds of 0.5 and 0.7. Transmission costs are calculated through Average MegaBytes (AM). In line with previous studies [48, 39], we compare methods in two configurations: (1) Synchronous, ignoring communication delays. (2) Asynchronous, simulating data transmission delay by retrieving roadside samples with the preceding timestamp.
@@ 4.2 Cooperative 3D Object Detection
Method M Sync. (AP@IoU = 0.5 / 0.7) Async. (AP@IoU = 0.5 / 0.7) AM
@@
No Fusion L 21.44 / 17.40 22.73 / 18.54 22.22 / 18.60 18.15 / 13.64 21.44 / 17.40 22.73 / 18.54 22.22 / 18.60 18.15 / 13.64 0
Late Fusion L 54.38 / 35.03 67.09 / 52.93 49.56 / 35.46 32.58 / 20.51 47.84 / 22.11 59.97 / 33.13 42.05 / 20.84 28.27 / 14.78 0.003
Early Fusion L 58.89 / 49.04 74.13 / 63.74 52.73 / 43.64 34.38 / 28.71 51.09 / 27.22 64.38 / 35.67 43.26 / 22.05 29.93 / 19.08 1.25
F-Cooper [2] L 60.27 / 45.33 74.91 / 61.87 56.64 / 38.95 35.85 / 21.32 55.56 / 32.03 70.11 / 44.81 50.01 / 23.64 32.48 / 17.31 0.25
CoAlign [17] L 66.42 / 55.56 70.81 / 63.27 66.75 / 54.15 58.36 / 43.01 58.43 / 35.07 63.32 / 39.61 56.64 / 31.83 50.65 / 31.00 0.25
HEAL [18] L 67.75 / 57.78 72.78 / 65.59 69.24 / 56.23 57.79 / 45.32 60.75 / 39.27 65.72 / 43.95 59.99 / 36.04 51.41 / 34.52 0.25
No Fusion C 3.45 / 2.61 5.08 / 1.56 0.69 / 0.29 0.22 / 0.10 3.45 / 2.61 5.08 / 1.56 0.69 / 0.29 0.22 / 0.10 0
Late Fusion C 12.82 / 7.56 27.33 / 16.53 2.38 / 1.93 1.87 / 0.92 11.08 / 4.68 23.88 / 10.19 1.94 / 1.14 1.59 / 0.58 0.003
V2X-ViT [36] C 17.51 / 10.54 32.28 / 19.93 3.42 / 2.64 2.34 / 1.39 15.61 / 7.24 29.19 / 13.58 2.95 / 1.71 2.05 / 0.92 0.25
CoAlign [17] C 19.33 / 12.67 37.01 / 21.05 3.89 / 2.73 2.45 / 1.41 17.33 / 8.61 33.41 / 14.11 3.27 / 1.74 2.17 / 0.97 0.25
HEAL [18] C 19.58 / 13.56 34.74 / 24.84 4.02 / 2.88 2.80 /1.53 17.42 / 9.53 31.23 / 17.43 3.43 / 1.96 2.47 / 0.98 0.25
No Fusion R 6.98 / 3.06 11.63 / 6.62 1.08 / 0.52 0.87 / 0.26 6.98 / 3.06 11.63 / 6.62 1.08 / 0.52 0.87 / 0.26 0
Late Fusion R 10.95 / 6.87 20.59 / 11.69 1.93 / 1.15 1.21 / 0.54 9.76 / 4.67 18.39 / 7.83 1.44 / 0.74 1.06 / 0.41 0.003
Early Fusion R 13.65 / 7.37 24.07 / 13.89 2.19 / 1.49 1.25 / 0.71 12.09 / 5.24 21.51 / 9.76 1.67 / 1.01 1.09 / 0.46 0.32
V2X-ViT [36] R 14.23 / 8.74 25.23 / 14.92 2.84 / 1.77 1.66 / 0.93 12.68 / 6.01 22.82 / 10.17 2.25 / 1.15 1.46 / 0.61 0.25
CoAlign [17] R 15.59 / 9.76 26.76 / 17.32 3.08 / 1.86 2.05 / 1.23 13.71 / 6.52 23.63 / 11.49 2.41 / 1.16 1.77 / 0.82 0.25
HEAL [18] R 15.82 / 10.08 27.02 / 18.24 3.56 / 1.91 1.83 / 1.12 13.97 / 7.56 24.11 / 12.97 2.85 / 1.23 1.62 / 0.75 0.25
5 Experiments
5.1 Implementation details
For the dataset split, single-agent 3D object detection, which includes roadside and vehicle-side scenarios, follows two distinct setups: (1) homogeneous split, where the dataset is divided into train/val/test sets containing 7000, 1500, and 1500 frames, respectively. The selection of samples in this setup is entirely random. (2) heterogeneous split, where the dataset is divided into train/val/test sets with 34, 8, and 8 sequences. The cooperative 3D object detection dataset is divided into train/val/test sets with 30, 5, and 5 sequences. Notably, all experimental results are assessed using the validation set. For the detection areas, the single-agent detection techniques cover a range of [0, 100] m in the x-direction and [-100, 100] m in the y-direction relative to the ego vehicle or roadside unit. In contrast, cooperative 3D object detection encompasses a broader area, extending from [-100, 100] m in both the x and y directions relative to the ego vehicle. For the ground truth, LiDAR or camera methods utilize all the annotations. Considering the limited coverage area of 4D Radar, 4D Radar-based methods only use labels within its field of view for training and evaluation. For the feature extractor, all cooperative detection methods depend on PointPillar [11] to extract BEV features from LiDAR or 4D Radar point clouds and use LSS [9] to extract BEV features from camera images.
@@ 5.1 Implementation details
Figure 6: Cooperative 3D object detection benchmarks on V2X-Radar-C under different transmission delay.
5.2 Benchmark results
Single-agent 3D Object Detection. The benchmark results for the V2X-Radar-I and V2X-Radar-V sub-datasets under a homogeneous split are detailed in Tab. 3 and Tab.  4. Further results under the heterogeneous split can be found in the supplementary materials. The experimental results in these tables clearly demonstrate that LiDAR-based methods achieve the highest performance. Although 4D radar-based methods manage a relatively sparse point cloud, they still outperform camera-based methods. Camera-based methods, restricted by their inability to utilize depth information, are less effective than LiDAR and 4D Radar-based methods.
Cooperative 3D Object Detection. Tab. 5 presents a quantitative comparison of several methods on V2X-Radar-C:
  • In comparison to the benchmark of single-vehicle perception, all methods involving cooperative perception exhibit a significant performance enhancement, highlighting its essential role in improving single-vehicle perception. Compared to the results from the Sync setting, the introduction of communication delay in Async settings led to a considerable performance decline. As shown in Tab. 5 and Fig. 6, when subjected to a 100ms transmission delay with a strict 0.7 IoU threshold, F-Cooper [2], CoAlign [17], and HEAL [18] using LiDAR point cloud experienced significant reductions of 13.30%, 20.49%, and 18.51% respectively. These findings underline the necessity of mitigating the effects of communication delay to ensure effective and robust cooperative perception.
@@ 6 Conclusion
In this paper, we introduce V2X-Radar, the first large-scale real-world multi-modal dataset featuring 4D Radar for cooperative perception, aiming to prompt cooperative perception with 4D Radar for a broader operational design domain (ODD). Our dataset focuses on challenging intersection scenarios and provides data collected under various times and weather conditions. It consists of 20K LiDAR frames, 40K camera images, and 20K 4D Radar data points, along with 350K annotated bounding boxes 5 categories. To support a wide range of perception research, we organise the dataset into three sub-datasets: V2X-Radar-C for cooperative perception, V2X-Radar-I for roadside perception, and V2X-Radar-V for single-vehicle perception. Furthermore, we have conducted extensive benchmarks of recent perception algorithms on these three sub-datasets.
Limitation and Future Work. Our dataset focuses on 3D object detection, and we simulate asynchronous cooperative perception with a fixed time delay. In the future, we plan to expand our V2X-Radar tasks to include tracking and trajectory prediction, and evaluate performance under real C-V2X communication delays.
Supplementary Material
@@
Table 6: Roadside 3D object detection benchmarks on V2X-Radar-I under the heterogeneous split. The vehicle category includes car, bus and truck. ‘M‘ means modality, and L, C, and R denote LiDAR, Camera, and 4D Radar, respectively.
Easy Moderate Hard Easy Moderate Hard Easy Moderate Hard
Pointpillars [11] L 66.19 / 86.11 61.28 / 78.82 61.28 / 78.82 33.32 / 49.80 27.79 / 41.28 27.79 / 41.28 64.87 / 74.11 56.55 / 64.95 56.55 / 64.95
SECOND [40] L 68.93 / 86.82 62.11 / 79.48 62.11 / 79.48 39.76 / 56.78 32.83 / 47.03 32.83 / 47.03 69.73 / 77.73 60.79 / 68.29 60.79 / 68.29
CenterPoint [47] L 70.87 / 87.62 65.62 / 80.22 65.63 / 80.22 57.66 / 71.58 47.29 / 59.34 47.29 / 59.34 73.08 / 78.15 65.46 / 70.27 65.46 / 70.27
PV-RCNN [21] L 82.08 / 92.27 74.75 / 84.81 74.75 / 84.81 67.33 / 80.39 56.72 / 66.69 56.72 / 66.69 76.59 / 82.04 67.11 / 72.51 67.11 / 72.51
SMOKE [15] C 3.04 / 16.87 2.99 / 16.64 2.99 / 16.63 0.05 / 0.09 0.04 / 0.07 0.04 / 0.07 1.12 / 8.25 1.07 / 7.47 1.07 / 7.39
BEVDepth [13] C 12.24 / 23.06 11.45 / 21.76 10.72 / 21.58 0.08 / 2.11 0.06 / 0.86 0.06 / 0.86 1.75 / 12.92 1.56 / 10.24 1.55 / 10.13
BEVHeight [42] C 12.94 / 24.41 11.61 / 23.55 11.40 / 23.35 1.01 / 3.11 0.08 / 1.29 0.08 / 1.24 3.31 / 18.14 2.01 / 12.24 2.01 / 12.22
BEVHeight++[41] C 14.07 / 25.12 13.57 / 24.81 13.25 / 24.62 1.39 / 4.87 0.13 / 1.63 0.13 / 1.62 4.20 / 19.69 2.57 / 12.81 2.57 / 12.78
RDIoU [20] R 4.31 / 25.74 3.85 / 24.24 3.85 / 24.24 1.55 / 8.64 1.16 / 6.96 1.16 / 6.96 1.94 / 10.91 1.55 / 9.65 1.55 / 9.65
RPFA-Net [33] R 4.49 / 28.39 4.18 / 26.39 4.18 / 26.39 1.76 / 7.82 1.32 / 7.04 1.32 / 7.04 2.36 / 12.61 1.95 / 10.94 1.95 / 10.94
Table 7: Single-vehicle 3D object detection benchmarks on V2X-Radar-V under the heterogeneous split. The vehicle category includes car, bus and truck. ‘M‘ means modality, and L, C, and R denote LiDAR, Camera, and 4D Radar, respectively.
Easy Moderate Hard Easy Moderate Hard Easy Moderate Hard
Pointpillars [11] L 72.03 / 80.43 64.65 / 72.87 64.65 / 71.87 39.01/44.17 33.37 / 40.17 33.37 / 40.17 72.46 / 75.99 61.97 / 65.82 61.97 / 65.82
SECOND [40] L 75.13 / 83.37 66.06 / 76.08 66.06 / 76.08 41.63 / 46.39 35.34 / 42.69 35.34 / 42.69 75.17 / 80.84 63.07 / 69.71 63.07 / 69.71
CenterPoint [47] L 77.67 / 85.22 67.24 / 78.26 67.24 / 78.26 53.77 / 57.67 48.08 / 54.68 48.08 / 54.68 83.92 / 89.51 72.69 / 75.69 72.69 / 75.69
PV-RCNN [21] L 83.34 / 87.91 77.37 / 81.24 77.37 / 81.24 62.33 / 63.18 56.87 / 59.87 56.87 / 59.87 86.68 / 91.54 76.73 / 78.05 76.73 / 78.05
SMOKE [15] C 0.31 / 7.28 0.16 / 3.81 0.16 / 3.74 0.03 / 0.18 0.02 / 0.11 0.02 / 0.11 0.08 / 1.89 0.07 / 1.25 0.07 / 1.25
BEVDepth [13] C 1.23 / 9.67 0.58 / 5 .57 0.58 / 5.41 0.06 / 1.02 0.04 / 0.71 0.04 / 0.71 1.01 / 3.74 0.05 / 1.91 0.05 / 1.91
BEVHeight [42] C 1.12 / 9.79 0.56 / 4.89 0.53 / 4.43 0.05 / 0.93 0.04 / 0.69 0.04 / 0.69 0.09 / 3.34 0.05 / 1.89 0.05 / 1.89
BEVHeight++[41] C 1.79 / 14.26 0.62 / 7.61 0.58 / 6.87 0.07 / 1.83 0.06 / 1.16 0.06 / 1.16 0.13 / 3.91 0.07 / 2.31 0.07 / 2.28
RDIoU [20] R 3.77 / 34.43 3.48 / 19.03 3.37 / 17.74 0.04 / 0.75 0.02 / 0.62 0.02 / 0.62 0.88 / 8.94 0.29 / 4.15 0.29 / 4.15
RPFA-Net [33] R 4.23 / 35.49 3.81 / 19.44 3.62 / 18.63 0.04 / 0.95 0.03 / 0.81 0.03 / 0.81 1.24 / 11.01 0.57 / 5.49 0.57 / 5.49
Figure 8: Diverse corner cases for single-vehicle autonomous driving. Each case is depicted by roadside multi-view images on the left and the vehicle-side image on the right. We use a red circle to mark the occluded objects.

### v3

@@ 4.1 Single-agent 3D Object Detection
Evaluation Metrics. The evaluation area covers [-100, 100] in x direction and [0, 100] in y direction from the ego-vehicle or roadside unit. Following the metrics in Rope3D [51], KITTI [6], we utilize the Average Precision (AP) at IoU thresholds of 0.5 and 0.7 as the evaluation metric.
@@
4.2 Cooperative 3D Object Detection
Evaluation Metrics. The evaluation area spans [-100, 100] metres in both the x and y directions relative to the ego vehicle. Similar to DAIR-V2X [54] and V2V4Real [40], we group various vehicle types into the same class and concentrate solely on vehicle detection. The performance of object detection is evaluated using Average Precision (AP) at IoU thresholds of 0.5 and 0.7. Transmission costs are calculated through Average MegaBytes (AM). In line with previous studies [54; 40], we compare methods in two configurations: (1) Synchronous, ignoring communication delays. (2) Asynchronous, simulating the delay by retrieving roadside samples with the preceding timestamp.
5 Experiments
5.1 Implementation Details
For the dataset split, single-agent 3D object detection, which includes roadside and vehicle-side scenarios. The dataset is divided into train/val/test sets containing 7000, 1500, and 1500 frames, respectively. The selection of samples in this setup is entirely random. The cooperative 3D object detection dataset is divided into train/val/test sets with 30, 5, and 5 sequences. Notably, all experimental results are assessed using the validation set. For the detection areas, the single-agent detection techniques cover a range of [0, 100] m in the x-direction and [-100, 100] m in the y-direction relative to the ego vehicle or roadside unit. In contrast, cooperative 3D object detection encompasses a broader area, extending from [-100, 100] m in both the x and y directions relative to the ego vehicle. For the ground truth, LiDAR or camera methods utilize all the annotations. Considering the limited coverage area of 4D Radar, 4D Radar-based methods only use labels within its field of view for training and evaluation. For the feature extractor, all cooperative detection methods depend on PointPillar [11] to extract BEV features from LiDAR or 4D Radar point clouds and use LSS [9] to extract BEV features from images. All experiments are conducted using 4 RTX-3090 GPUs.
5.2 Benchmark Results
Single-agent 3D Object Detection. The benchmark results for the V2X-Radar-I and V2X-Radar-V sub-datasets under a homogeneous split are detailed in Tab. 3 and Tab.  4. The experimental results in these tables clearly demonstrate that LiDAR-based methods achieve the highest performance. Although 4D radar-based methods manage a relatively sparse point cloud, they still outperform camera-based methods. Camera-based methods, restricted by their inability to utilize depth information, are less effective than LiDAR and 4D Radar-based methods.
Table 3: Roadside 3D object detection benchmarks on V2X-Radar-I under homogeneous split. The vehicle category includes car, bus and truck. "M" means modality, and "L", "C", and "R" denote LiDAR, Camera, and 4D Radar, respectively.
@@ 5.2 Benchmark Results
Table 4: Single-vehicle 3D object detection benchmarks on V2X-Radar-V under homogeneous split. The vehicle category includes car, bus and truck. "M" means modality, and "L", "C", and "R" denote LiDAR, Camera, and 4D Radar, respectively.
@@
Cooperative 3D Object Detection. A quantitative comparison of representative methods on the V2X-Radar-C sub-dataset is shown in Tab. 5, and the main findings can be summarized as follows:
  • In comparison to the benchmark of single-vehicle perception, all methods involving cooperative perception exhibit a significant performance enhancement, highlighting its essential role in improving single-vehicle perception.
  • Compared to the results from the Sync setting, the introduction of communication delay in Async settings led to a considerable performance decline. As shown in Tab. 5 and Fig. 7, when subjected to transmission delay with a strict 0.7 IoU threshold, F-Cooper [2], CoAlign [18], and HEAL [17] using LiDAR point cloud experienced significant reductions of 13.30%, 20.49%, and 18.51% respectively. These findings underline the necessity of mitigating the effects of communication delay to ensure effective and robust cooperative perception.
@@
Method M Sync. (AP@IoU = 0.7 / 0.5) Async. (AP@IoU = 0.7 / 0.5) AM
@@
No Fusion L 17.40 / 21.44 18.54 / 22.73 18.60 / 22.22 13.64 / 18.15 17.40 / 21.44 18.54 / 22.73 18.60 / 22.22 13.64 / 18.15 0
Late Fusion L 35.03 / 54.38 52.93 / 67.09 35.46 / 49.56 20.51 / 32.58 22.11 / 47.84 33.13 / 59.97 20.84 / 42.05 14.78 / 28.27 0.003
Early Fusion L 49.04 / 58.89 63.74 / 74.13 43.64 / 52.73 28.71 / 34.38 27.22 / 51.09 35.67 / 64.38 22.05 / 43.26 19.08 / 29.93 1.25
F-Cooper [2] L 45.33 / 60.27 61.87 / 74.91 38.95 / 56.64 21.32 / 35.85 32.03 / 55.56 44.81 / 70.11 23.64 / 50.01 17.31 / 32.48 0.25
CoAlign [18] L 55.56 / 66.42 63.27 / 70.81 54.15 / 66.75 43.01 / 58.36 35.07 / 58.43 39.61 / 63.32 31.83 / 56.64 31.00 / 50.65 0.25
HEAL [17] L 57.78 / 67.75 65.59 / 72.78 56.23 / 69.24 45.32 / 57.79 39.27 / 60.75 43.95 / 65.72 36.04 / 59.99 34.52 / 51.41 0.25
No Fusion C 2.61 / 3.45 1.56 / 5.08 0.29 / 0.69 0.10 / 0.22 2.61 / 3.45 1.56 / 5.08 0.29 / 0.69 0.10 / 0.22 0
Late Fusion C 7.56 / 12.82 16.53 / 27.33 1.93 / 2.38 0.92 / 1.87 4.68 / 11.08 10.19 / 23.88 1.14 / 1.94 0.58 / 1.59 0.003
V2X-ViT [41] C 10.54 / 17.51 19.93 / 32.28 2.64 / 3.42 1.39 / 2.34 7.24 / 15.61 13.58 / 29.19 1.71 / 2.95 0.92 / 2.05 0.25
CoAlign [18] C 12.67 / 19.33 21.05 / 37.01 2.73 / 3.89 1.41 / 2.45 8.61 / 17.33 14.11 / 33.41 1.74 / 3.27 0.97 / 2.17 0.25
HEAL [17] C 13.56 / 19.58 24.84 / 34.74 2.88 / 4.02 1.53 / 2.80 9.53 / 17.42 17.43 / 31.23 1.96 / 3.43 0.98 / 2.47 0.25
No Fusion R 3.06 / 6.98 6.62 / 11.63 0.52 / 1.08 0.26 / 0.87 3.06 / 6.98 6.62 / 11.63 0.52 / 1.08 0.26 / 0.87 0
Late Fusion R 6.87 / 10.95 11.69 / 20.59 1.15 / 1.93 0.54 / 1.21 4.67 / 9.76 7.83 / 18.39 0.74 / 1.44 0.41 / 1.06 0.003
Early Fusion R 7.37 / 13.65 13.89 / 24.07 1.49 / 2.19 0.71 / 1.25 5.24 / 12.09 9.76 / 21.51 1.01 / 1.67 0.46 / 1.09 0.32
V2X-ViT [41] R 8.74 / 14.23 14.92 / 25.23 1.77 / 2.84 0.93 / 1.66 6.01 / 12.68 10.17 / 22.82 1.15 / 2.25 0.61 / 1.46 0.25
CoAlign [18] R 9.76 / 15.59 17.32 / 26.76 1.86 / 3.08 1.23 / 2.05 6.52 / 13.71 11.49 / 23.63 1.16 / 2.41 0.82 / 1.77 0.25
HEAL [17] R 10.08 / 15.82 18.24 / 27.02 1.91 / 3.56 1.12 / 1.83 7.56 / 13.97 12.97 / 24.11 1.23 / 2.85 0.75 / 1.62 0.25
@@
Figure 7: Cooperative 3D object detection benchmarks on V2X-Radar-C under different transmission delay. Each data point reflects the average three separate experiments.
Adverse-weather Robustness through 4D Radar. We conducted ablation studies on a subset of vehicle-side frames collected under adverse weather conditions (rain, fog, and snow), comparing LiDAR-only, 4D Radar-only, and LiDAR-4D Radar fusion models, as shown in Tab. 6. The results are as summarized as follows:
  • While generally lagging behind LiDAR-only method under normal conditions, 4D Radar-only model outperforms LiDAR-only model by approximately 1–2% mAP in adverse weather conditions, highlighting its resilience to environmental degradation.
  • The LiDAR-4D Radar fusion model consistently achieves the highest accuracy across both LiDAR-only and 4D Radar-only models, demonstrating the complementary sensing strengths of the two modalities, where LiDAR provides precise spatial geometry while 4D Radar contributes robustness under adverse conditions.
Table 6: Ablation results under adverse weather conditions (rain, heavy fog, and snow). 4D Radar-only method outperforms LiDAR-only model by 1-2% mAP in harsh conditions, and the fusion model consistently achieves the best results, demonstrating their complementary strengths.
Method Modality Vehicle (IoU = 0.5) Pedestrian (IoU = 0.25) Cyclist (IoU = 0.25)
Easy Moderate Hard Easy Moderate Hard Easy Moderate Hard
Pointpillars [11] LiDAR 47.23 43.12 42.15 21.12 17.52 16.87 29.32 25.34 24.81
4D Radar 48.35 44.80 43.91 20.14 16.54 16.20 27.85 23.66 22.68
M2-Fusion [52] LiDAR + 4D Radar 53.61 50.18 49.72 25.11 20.53 19.59 33.42 28.24 27.12
@@ 6 Conclusion
In this work, we introduced V2X-Radar, the first large-scale real-world multi-modal dataset featuring 4D Radar for cooperative perception. Our dataset focuses on challenging intersection scenarios and provides data collected under diverse times and weather conditions. Beyond releasing a dataset and benchmarks, our study revealed two key insights for the research community: (i) Severe performance degradation under asynchronous communication settings. This finding exposes a critical weakness in the delay robustness of current cooperative perception methods. (ii) The distinctive advantage of 4D Radar, which delivers reliable perception in adverse weather and serves as a valuable complement to LiDAR- and camera-based approaches. By releasing V2X-Radar, we not only fill the 4D Radar gap in cooperative perception research but also offer a platform for studying these challenges and validating solutions. We hope this dataset and benchmark will spark future work on delay-tolerant cooperative perception models, robust cross-modal fusion strategies.
Broader Impacts. Although the proposed benchmark covers various driving scenes, due to differences in sensor configurations, models trained on this dataset may not generalize well to other vehicle-side or roadside unit platforms, thus failing to ensure the safety of autonomous driving.
7 Limitation and Future Work
Our current dataset primarily focuses on 3D object detection, providing a foundation for studying cooperative perception under diverse sensing modalities and scenarios. However, it is limited in temporal coverage and task diversity, as it does not yet include sequential or predictive perception tasks. In future, we plan to extend the V2X-Radar dataset toward more comprehensive cooperative perception benchmarks. We will introduce multi-object tracking and trajectory prediction to capture temporal dynamics and enable spatiotemporal reasoning, and develop an Occupancy Prediction task formulated as voxel-level semantic segmentation with both dynamic and static classes. These extensions aim to enable richer scene understanding and advance V2X-Radar toward a complete world-model-based perception benchmark.
Appendix A Summary
This supplementary document is organized as follows:
  • Appendix B provides detailed descriptions of the time synchronization mechanisms adopted in V2X-Radar, including synchronization within and across vehicle-side and roadside platforms.
  • Appendix C presents the robustness evaluation of time synchronization under varying transmission delays, analyzing the impact of temporal asynchrony on cooperative perception performance.
  • Appendix D compares the V2X-Radar single-agent subsets with representative autonomous driving datasets, highlighting their advantages in 4D Radar sensing, adverse-weather coverage, and multi-pass data collection.
  • Appendix E introduces quantitative calibration metrics for LiDAR-Camera and LiDAR-4D Radar pairs, providing reprojection and alignment error analyses that demonstrate the dataset’s high calibration precision.
  • Appendix F showcases visual examples of diverse data collection scenarios, including corner cases for single-vehicle perception and variations across different times of day and weather conditions.
Appendix B Time Synchronization Details
Time synchronization enables the simultaneous sampling of data from the same scene by various sensors on both the vehicle platform and roadside units. For a collaborative perception dataset, achieving precise time alignment is critical not only among the diverse sensors within a single platform (be it the vehicle or the roadside unit) but also across sensors spanning both platforms.
@@ Appendix B Time Synchronization Details
Appendix C Robustness Evaluation of Time Synchronization
To comprehensively evaluate the robustness of existing cooperative perception methods under temporal misalignment, we adopt the time synchronization protocol used in DAIR-V2X [54] and V2V4Real [40]. Specifically, we simulate transmission delays of 0 ms (synchronous), 100 ms, 200 ms, 300 ms, and 400 ms between vehicle-side and roadside sensors on the V2X-Radar-C dataset to analyze the sensitivity of perception performance to temporal offsets. The results, summarized in Tab. 7, demonstrate a consistent degradation trend as the delay increases, highlighting the detrimental impact of temporal asynchrony on cooperative perception.
Table 7: Evaluation of the time synchronization robustness on V2X-Radar-C dataset under varying transmission delays. "M" denotes the modality ("L": LiDAR, "C": Camera, "R": 4D Radar). "AM" indicates the average transmission cost in MB.
[…]
No Fusion L 21.44 21.44 21.44 21.44 21.44 0.00
Late Fusion L 54.38 47.84 42.91 40.62 38.66 0.003
F-Cooper [2] L 60.27 55.56 48.10 45.68 44.30 1.25
CoAlign [18] L 66.42 58.43 54.60 51.54 49.82 0.25
HEAL [17] L 67.75 60.75 56.57 53.12 52.17 0.25
No Fusion C 3.45 3.45 3.45 3.45 3.45 0.00
Late Fusion C 12.82 11.08 10.00 9.58 9.12 0.003
F-Cooper [2] C 17.51 15.61 14.01 13.27 12.87 1.25
CoAlign [18] C 19.33 17.33 16.04 15.00 14.50 0.25
HEAL [17] C 19.58 17.42 16.45 15.35 15.08 0.25
No Fusion R 6.98 6.98 6.98 6.98 6.98 0.00
Late Fusion R 10.95 9.76 8.98 8.18 7.79 0.003
F-Cooper [2] R 14.23 12.68 11.81 10.79 10.12 1.25
CoAlign [18] R 15.59 13.71 13.10 12.10 11.08 0.25
HEAL [17] R 15.82 13.97 13.45 12.40 11.25 0.25
Appendix D Dataset Comparison from a Single-agent Perspective
Although V2X-Radar was originally built for cooperative perception, it also includes two single-agent subsets, V2X-Radar-I and V2X-Radar-V, that have unique strengths. They feature real 4D Radar sensing, broad coverage of challenging weather conditions, and repeated multi-pass recordings that enable accurate scene reconstruction and HD map generation. To highlight these strengths, we present a comparison table (Tab. 8) that focuses on single-agent datasets. Our subsets are compared with well-known autonomous driving datasets such as BDD100K [53], KITTI [6], nuScenes [1], Waymo [29], Rope3D [51], VOD [21], and K-Radar [20]. The results show that, while V2X-Radar was designed for cooperative use, its single-agent subsets provide higher-quality 4D Radar data, better performance in adverse weather, and richer multi-pass coverage, making them a strong resource for research on robust perception, scene reconstruction, and mapping.
Table 8: Comparison of our single-agent dataset with existing autonomous driving datasets. Although V2X-Radar is designed for cooperative perception, its single-agent subsets provide unique advantages in real 4D Radar sensing, adverse-weather coverage, and multi-pass data for scene reconstruction and HD mapping.
Dataset Location Num. Data LiDAR Camera 4D Radar GPS / RTK Day & Night Adverse Weather Multi-pass Coverage
BDD100K [53] USA 100K
KITTI [6] DE 15K
nuScenes [1] Singapore 40K
Waymo [29] USA 230K
Rope3D [51] China 40K
VOD [21] Netherlands 8.7K
K-Radar [20] South Korea 35K
V2X-Radar-I China 20K
V2X-Radar-V China 20K
Appendix E Quantitative Calibration Metrics
To quantitatively assess sensor calibration accuracy, we evaluate both LiDAR-Camera and LiDAR-4D Radar calibration errors.
For LiDAR-Camera calibration, the reprojection error (RPE) is used as the evaluation metric. A checkerboard target is employed to extract 3D corner points in the LiDAR frame and their corresponding pixel coordinates in the camera image. Using the estimated calibration parameters, the 3D points are projected onto the image plane, and the mean pixel-wise L2 distance between the projected and true corner locations is reported as the RPE.
For LiDAR-4D Radar calibration, the alignment error (AE) is used. Corner reflectors are employed to obtain paired 3D points in both LiDAR and Radar frames. Radar points are transformed into the LiDAR coordinate system based on the estimated extrinsics, and the mean Euclidean (L2) distance between the transformed Radar points and their corresponding LiDAR points is computed as the AE.
As summarized in Tab. 9, the V2X-Radar dataset achieves LiDAR-Camera RPEs of 0.26-0.33 pixels (well below the acceptable 0.5 pixel threshold) and LiDAR-4D Radar AEs of 2.23-2.35 cm (within the acceptable < 5 cm range). These results demonstrate the dataset’s high calibration precision and cross-sensor consistency across both vehicle-side and roadside platforms.
Table 9: Calibration accuracy across different sensor pairs on both vehicle-side and roadside platforms. LiDAR-Camera calibration is evaluated by the re-projection error (RPE, pixel), and LiDAR-4D Radar calibration by the alignment error (AE, cm). Lower values indicate more accurate extrinsic calibration.
[…]
Vehicle-side LiDAR – CAM 0.33 /
LiDAR – 4D Radar / 2.35
Roadside LiDAR – CAM_LEFT 0.26 /
LiDAR – CAM_FRONT 0.27 /
LiDAR – CAM_RIGHT 0.30 /
LiDAR – 4D Radar / 2.23
Appendix F Data Collection Scenarios
Our data collection covers a wide range of weather conditions and times of day, with a particular focus on challenging corner cases for single-vehicle autonomous driving.
Figure 9 further depicts the diversity of the collection scenarios, spanning multiple times of day and varying meteorological conditions As shown in Figure 9(a), the temporal segments include morning, afternoon, dusk, and nighttime scenes. Figure 9(b) demonstrates the range of weather conditions, from clear and sunny to fog, rain, and snow.